A round body measures 5 kg and has a radius of 30 cm. What moment of force should be exerted on it to increase the frequency from 5 rpm to 10 rpm, in 4 seconds?


1
Expert's answer
2021-04-02T10:34:50-0400
"\tau=I\alpha=\frac{2}{5}mr^2\\\tau=\frac{2}{5}(5)(0.3)^2\frac{2\pi}{60}\frac{10-5}{4}=0.0236\ Nm"